# OpenGraph: Towards Open Graph Foundation Models

### v2

@@ 4.2. Overall Performance Comparison (RQ1)
We begin by comparing the zero-shot performance of OpenGraph with the few-shot performance of the baselines in both link prediction and node classification. The results are displayed in Table 1. Based on these results, we have made the following observations:
@@
O1: Predominant Performance of OpenGraph. Our model demonstrates superior performance across all 8 datasets in various categories. Notably, this advantage is achieved solely through pre-training on a generated dataset, without any overlap with the downstream graph data. This highlights the remarkable ability of our model to generalize across different datasets. We attribute this advantage to three key factors: i) The unified graph tokenizer, which effectively bridges the distribution gap between the pre-training and target datasets. ii) The scalable graph transformer, which excels at learning relations and capturing important structural features, resulting in high-quality graph representations. iii) The effective pre-training with LLM-generated graph data, which equips our model with versatile forecasting abilities for unseen downstream datasets. Overall, these design choices contribute to the outstanding generalization capabilities of our model across diverse datasets.
@@
O2: Limitations of Existing Pretraining Methods. Observations reveal that pre-training techniques, such as GraphCL and DGI, do not consistently surpass their foundational models, like GIN and GCN, which were trained solely on few-shot data. This pattern suggests that the pre-training methods may hinder rather than help performance when models are applied to different datasets. Such underperformance is likely due to a substantial shift in data distribution between the pre-training and target datasets, which can lead the model to become overfitted to the pre-training data, thus impairing its ability to adapt to new graph structures in the downstream tasks. Furthermore, prompt tuning methods like GraphPrompt and GPF sometimes degrade more severely compared to full-model fine-tuning approaches, which indicates an even greater susceptibility to overfitting to distinctive patterns within the pre-training dataset.
@@
O3: Tackling Node Classification by Structure Learning. Our OpenGraph exhibits a marked improvement in node classification tasks, which underscores the efficacy of employing our pre-trained link predictor to discern the connections between ordinary nodes and special nodes representing classes. This advantage hinges on OpenGraph’s robust capacity to generalize, as it successfully transfers knowledge both across different datasets and between distinct tasks. The credit for this versatility goes to the universal applicability and flexibility of our graph tokenization and graph encoding.
4.3. Investigation on Graph Tokenizer (RQ2)
@@
In this section, we study the effectiveness of our unified graph tokenization module. This investigation involves an evaluation of the impact of the smoothed adjacency matrix and a comparison of our topology-aware projection method with alternative compression methods. The evaluation results are presented in Figure 2, based on which we make the following analysis.
@@
Superiority of topology-aware projection. To assess the effectiveness of our topology-aware projection based on fast SVD, we compare it to three alternative projection modules (see Appendix A.2.5 for more details). The results are presented in Figure 2(c) and 2(d). We make the following observations:
@@
  • One-hot encoding. This graph projection strategy learns id-corresponding embeddings across datasets. It performs poorly in our zero-shot evaluation, highlighting the difficulty of transferring dataset-specific parameters such as node embeddings to unseen datasets that lack overlapping node tokens.
@@
  • Degree embeddings. This method learns degree-specific embeddings. It demonstrates a significant performance disadvantage compared to our projection scheme. This can be attributed to the substantial semantic gap that exists for the same degree number across different graphs. Additionally, this approach oversimplifies the topology features by considering only the number of direct links, which limits its ability to capture more nuanced structural patterns and can be detrimental to the graph projection.
@@
  • Random projection. It randomly assigns an unlearnable embedding vector to each node. The results demonstrates its advantages over the other two variants. However, due to the low representation efficiency of its uniform distribution, this method also exhibits poorer performance compared to our method.
Refer to captionRefer to caption
(a) OGBL-ddi dataset.
Refer to captionRefer to caption
(b) Movielens-1M dataset.
Refer to captionRefer to caption
(c) OGBL-ddi dataset.
Refer to captionRefer to caption
(d) Movielens-1M dataset.
Figure 2. Influence of graph tokenizer configurations.
4.4. Influence of Pre-training Datasets (RQ3)
@@
To evaluate the effectiveness of our knowledge distillation from the LLM, we compare the performance of OpenGraph networks that are pre-trained with different datasets. In our experiments, we employ three ablated versions of our graph generation algorithm, namely -Norm, -Loc, and -Topo (further details below). Additionally, we incorporate two real datasets, Yelp2018 and Gowalla, for pre-training, which are unrelated to the test datasets used in this experiment. Moreover, we include the ML-10M dataset, which is related to both ML-1M and itself, in the test datasets. The evaluation results are summarized in Table 2. We draw the following conclusions:
@@
Superiority of our generated data. Our generated dataset, referred to as Gen, achieves the best performance on all test datasets except for ML-1M and ML-10M. Notably, ML-10M, which is closely related to these two datasets, achieves the best performance in these cases. This finding highlights the superior generalization ability of our generated datasets, which equips the OpenGraph model with the capability of universal topological structure learning.
@@
Effectiveness of individual generation techniques. We conduct ablation study by removing the dynamic probability normalization (-Norm), locality incorporation (-Loc), and graph topological pattern injection (-Topo) modules. The removal of these modules from the graph generation algorithm leads to a significant drop in performance for the downstream prediction. This demonstrates the effectiveness of these three key modules for our generated graphs.
@@
Real-world data can be misleading. Using real-world datasets such as Yelp2018 and Gowalla does not yield transferable graph learning capabilities that provide an advantage over our Gen dataset. This highlights the limitation of relying solely on insufficient or biased real-world data for cross-dataset pre-training.
@@
Pretraining on related datasets is useful. Models pre-trained on ML-10M exhibit superior performance on ML-1M and ML-10M, which are directly related datasets. This indicates that dataset-wise similarity can aid in the cross-dataset knowledge transferring.
Table 2. Influence of utilizing different pre-training datasets.
@@
Dataset -Norm -Loc -Topo Yelp2018 Gowalla ML-10M Gen
ogbl-ddi 0.0737 0.0893 0.0656 0.0588 0.0770 0.0692 0.0921
ML-1M 0.0572 0.1680 0.0850 0.0599 0.0485 0.2030 0.1911
ML-10M 0.0982 0.1636 0.1017 0.1629 0.0910 0.2698 0.2370
Cora 0.4985 0.4864 0.4342 0.3715 0.5943 0.2780 0.7504
Citeseer 0.3944 0.3691 0.5743 0.2651 0.4300 0.2003 0.6097
Pubmed 0.4501 0.5015 0.4876 0.3317 0.5148 0.3652 0.6869
4.5. Impact of Sampling in Transformer (RQ4)
We examine the influence of token sequence sampling and anchor sampling in our scalable graph transformer architecture. The metrics for efficiency and performance are summarized in Table 3. Our evaluation focuses on the GPU memory costs and the running time during both the training and testing processes. The OGBL-ddi and ML-10M datasets are used for evaluation. Additionally, we assess the model performance after end-to-end training on these datasets. The following observations are made for the ablated versions.
  • -Seq-Anc: This variant eliminates both token sequence sampling and anchor sampling, leading to out-of-memory (OOM) errors when applied to the larger ML-10M dataset. In the evaluation on OGBL-ddi, it exhibits the lowest memory and time efficiency, while its performance is inferior to our OpenGraph model.
  • -Anc: This version incorporates only sequence sampling into the graph transformer. It significantly reduces memory costs during the training phase. Additionally, by focusing on the current training context, it achieves the best performance on OGBL-ddi.
  • -Seq: This model removes token sequence sampling from our OpenGraph architecture, resulting in a significant decrease in training efficiency. However, the anchor sampling strategy in this version greatly reduces computational costs during the test phase. Despite the computational benefits, the anchor sampling strategy leads to a drop in performance compared to our OpenGraph.
Table 3. Impact of sampling strategies on the efficiency and model performance in the scalable graph transformer.
OGBL-ddi Train Mem Test Mem Train Time Test Time Recall@20
-Seq-Anc 5420MiB 1456MiB 22.72s 13.88s 0.0966
@@ 4.5. Impact of Sampling in Transformer (RQ4)
OpenGraph 2358MiB 1202MiB 15.45s 12.09s 0.1006
ML-10M Train Mem Test Mem Train Time Test Time Recall@20
-Seq-Anc OOM OOM
@@
OpenGraph 4470MiB 4550MiB 68.79s 54.17s 0.2816
4.6. Impact of Model Scale (RQ5)
@@
Hidden dimensionality. In contrast to the number of graph transformer layers, the hidden dimensionality leads to quadratic growth in computational time. This reflects the rapid increase in model capacity required to accommodate more complex structural data. Consequently, we observe significant improvements in model performance, surpassing the performance curve for the graph transformer layers. Despite the increased model capacity, the growth in hidden dimensionality facilitates the $\mathbb{R}^{N}\rightarrow\mathbb{R}^{d}$ projection, reducing information loss and enhancing the quality of the graph token sequence for the subsequent graph transformer layers.
Refer to captionRefer to caption
@@
Figure 3. The impact of OpenGraph model scale on downstream data performance and training/testing time (seconds).
4.7. End-to-End Training Performance (RQ6)
@@
5. Related Work
Graph Neural Networks. Graph Neural Networks (GNNs) have gained significant attention in recent years due to their ability to effectively capture and model complex relationships within graph-structured data (wu2020comprehensive, ; zhang2022graph, ; chen2020simple, ). GNNs utilize message passing schemes to propagate information across the graph and aggregate information from the neighboring nodes for node representations (jin2021universal, ; yuan2020xgnn, ). Several representative graph neural architectures have been developed. For example, Graph Convolutional Networks (GCNs) adapt the concept of convolution to operate on graph-structured data and enable nodes to exchange messages with their neighbors (gao2018large, ; zhang2021lorentzian, ). Additionally, Graph Attention Networks (GANs) are introduced to address the limitations of traditional GCNs in capturing complex relationships and varying importance levels between nodes in a graph (zhang2022graph, ; liao2019efficient, ). The backbone of OpenGraph draws inspiration from the powerful Graph Transformer (yun2019graph, ; hu2020heterogeneous, ) in capturing long-range dependencies and global information in graphs.
@@ 5. Related Work
Self-Supervised Learning on Graphs. To tackle the issue of limited labeled data in graph tasks, self-supervised graph learning techniques, such as contrastive and generative approaches, provide promising solutions (wu2021selftkde, ; lee2022augmentation, ; xia2023automated, ). These methods leverage the inherent structure and patterns within graphs to train models in a data-efficient manner (xiao2022decoupled, ; yang2023graph, ). Specifically, graph contrastive learning frameworks focus on creating meaningful representations by contrasting positive and negative samples. Notable models in this area include GraphCL (you2020graph, ) and SGL (wu2021self, ), which utilize stochastic data augmenters. Additionally, adaptive augmentation schemes such as JOAO (you2021graph, ) and GCA (zhu2021graph, ) have been proposed. Another paradigms, DGCL (li2021disentangled, ) and UMGRL (mo2023disentangled, ), addresses the challenge of disentangling factors within contrastive learning. However, existing solutions often struggle to generalize well across diverse downstream graph scenarios due to their strong assumptions of a consistent graph token set. This work proposes OpenGraph, which enhances graph model generalization across different graph tasks.
@@
LLM-based Methods for Graph Analysis. Recent advancements in the field of large language models (LLMs) have sparked researchers’ interest in leveraging their potential to enhance graph comprehension and analysis (chai2023graphllm, ; fatemi2023talk, ; ren2023representation, ; guo2024graphedit, ). Notably, GraphLLM (chai2023graphllm, ) and GraphQA (fatemi2023talk, ) are notable examples of efforts aimed at transforming graphs into natural language descriptions. Leveraging such descriptions as inputs for LLMs enables improved interpretation and reasoning on graph data. Additionally, techniques like instruction tuning in GraphEdit (guo2024graphedit, ) and GraphGPT (tang2023graphgpt, ) facilitate fine-tuning LLMs by incorporating rich textual information from text-attributed graphs. However, the success of these methods heavily relies on the availability of high-quality textual features associated with graph nodes. Acquiring such features can be challenging in certain graph domains. For example, privacy concerns can hinder obtaining high-quality text-attributed graphs in user behavior graphs, and constructing comprehensive textual descriptions for neurons in a brain’s neuronal graph can be difficult. Therefore, there is a need to develop a graph model that can capture universal structural patterns from graphs, even in the absence of textual data.
6. Conclusion
@@
The main focus of this research is to develop a highly adaptable framework capable of accurately capturing and comprehending the intricate topological patterns found in diverse graph structures. By harnessing the potential of the proposed model, our intention is to significantly improve the model’s ability to generalize and perform effectively in zero-shot graph learning tasks encompassing a diverse spectrum of downstream applications. To further improve the efficiency and robustness of OpenGraph, we build our model upon a scalable graph transformer architecture together with a LLM-enhanced data augmentation mechanism. Through extensive experiments conducted on multiple benchmark datasets, we have demonstrated the exceptional generalization abilities of our model. This study makes an initial attempt to exploit the graph fundation model. In future work, our plan is to empower our framework with the capability to automatically discover noisy connections and influential structures with counterfactual learning, while learning the universal and transferable structural patterns of diverse graphs.

### v4

@@ 4.2 Overall Performance Comparison (RQ1)
Comparing the zero-shot performance of OpenGraph with the few-shot performance of baselines in link prediction and node classification (Table 1), we have the following observations:
@@
Predominant Performance of OpenGraph. Our model outperforms baselines on all 8 datasets in different categories without using any overlapping data between pre-training and downstream tasks, showcasing its remarkable ability to generalize. This advantage can be attributed to three key factors: i) the unified graph tokenizer, bridging the gap between pre-training and target datasets; ii) the scalable graph transformer, capturing important structural features and learning relations effectively; and iii) effective pre-training with LLM-generated graph data, equipping our model with versatile forecasting abilities. Overall, these design choices contribute to our model’s outstanding generalization capabilities across diverse datasets.
@@
Limitations of Existing Pretraining Methods. Pre-training methods like GraphCL and DGI do not consistently outperform foundational models (e.g., GIN and GCN) trained on few-shot data, suggesting that they may hinder performance when applied to different datasets. This is due to a significant shift in data distribution between pre-training and target datasets, leading to overfitting and impairing adaptability to new graph structures in downstream tasks. Prompt tuning methods like GraphPrompt and GPF sometimes degrade more severely compared to full-model fine-tuning approaches, indicating a higher susceptibility to overfitting to distinctive patterns within the training set.
@@
Tackling Node Classification by Structure Learning. Our OpenGraph shows significant improvement in node classification tasks, highlighting the effectiveness of using our pre-trained link predictor to identify connections between ordinary nodes and special nodes representing classes. This advantage relies on OpenGraph’s strong ability to generalize, transferring knowledge across datasets and tasks. The credit for this versatility goes to the universal applicability and flexibility of our graph tokenization and encoding.
4.3 Investigation on Graph Tokenizer (RQ2)
@@
In this section, we study the effectiveness of our graph tokenizater by evaluating the impact of the smoothed adjacency matrix and comparing our projection method with alternative compression methods. The results are summarized in Figure 2.
@@
Superiority of topology-aware projection. To assess the effectiveness of our topology-aware projection based on SVD, we compare it to three alternative projection methods (see Appendix A.3.5 for details). The results are presented in Figure 2(c) and 2(d). We make the following observations:
@@
  • One-hot encoding. This approach learns id-corresponding embeddings across datasets. It performs poorly in the zero-shot evaluation, highlighting the difficulty of transferring dataset-specific parameters like node embeddings to unseen datasets that lack overlapping node tokens.
@@
  • Degree embeddings. This method learns degree-specific embeddings. It performs significantly worse than our projection scheme. This is because there is a substantial semantic gap for the same degree number across different graphs. Moreover, it oversimplifies topology features by considering only the number of direct links, limiting its ability to capture nuanced structural patterns and adversely affecting graph projection.
@@
  • Random projection. It randomly assigns unlearnable embedding vectors to nodes. It outperforms the other two variants, but its performance is still inferior to our method due to the low representation efficiency of its uniform distribution.
4.4 Influence of Pre-training Datasets (RQ3)
@@
To evaluate the effectiveness of our knowledge distillation from the LLM, we compare the performance of OpenGraph networks pre-trained with different datasets. We use three ablated versions of our graph generation algorithm, namely -Norm, -Loc, and -Topo. Additionally, we incorporate two real datasets, Yelp2018 and Gowalla, for pre-training, which are unrelated to the test datasets. The ML-10M dataset, related to ML-1M and itself, is also included. The evaluation results are summarized in Table 2. We draw the following conclusions:
(a) Recall@20
(b) Recall@40
(c) Recall@20
(d) Recall@40
Figure 2: Influence of graph tokenizer configurations.
@@
Superiority of our generated data. Our generated dataset (Gen) achieves the best performance on all test datasets except for ML-1M and ML-10M. Notably, ML-10M, which is closely related to these two datasets, achieves the best performance in these cases. This finding highlights the superior generalization ability of our generated datasets, which equips the OpenGraph model with the capability of universal topological structure learning.
@@
Impact of individual generation techniques. We conduct ablation study by removing the dynamic probability normalization (-Norm), locality incorporation (-Loc), and graph topological pattern injection (-Topo) modules. The removal of these modules from the generation algorithm leads to a significant drop in performance for the downstream prediction. This demonstrates the importance of these key modules for our generated graphs.
@@
Real-world data can be misleading. Using real-world datasets (Yelp and Gowalla) does not yield transferable graph learning capabilities that provide an advantage over our Gen data. This highlights the limitation of relying solely on insufficient or biased real-world data for cross-data pre-training.
@@
Pre-training on related datasets is useful. Models pre-trained on the ML-10M dataset exhibit superior performance when testing on ML-1M and ML-10M, which are directly related datasets. This indicates that dataset-wise similarity can aid in the cross-dataset knowledge transferring.
Table 2: Impact of using different pre-training datasets.
@@
Data -Norm -Loc -Topo Yelp Gowalla ML10M Gen
ogbl-ddi .0737 .0893 .0656 .0588 .0770 .0692 .0921
ML1M .0572 .1680 .0850 .0599 .0485 .2030 .1911
ML10M .0982 .1636 .1017 .1629 .0910 .2698 .2370
Cora .4985 .4864 .4342 .3715 .5943 .2780 .7504
Citeseer .3944 .3691 .5743 .2651 .4300 .2003 .7221
Pubmed .4501 .5015 .4876 .3317 .5148 .3652 .6869
4.5 Impact of Sampling in Transformer (RQ4)
We examine the influence of token sequence sampling and anchor sampling in our scalable graph transformer architecture. The metrics for efficiency and performance are summarized in Table 3. Our evaluation focuses on the GPU memory costs and the running time during both the training and testing processes. Additionally, we assess the model performance after end-to-end training. The following observations are made for the ablated versions.
  • -S-A: It eliminates both token sequence sampling and anchor sampling, leading to out-of-memory (OOM) errors when applied to the larger ML-10M data. When evaluated on OGBL-ddi, it exhibits the lowest memory and time efficiency, while its performance is inferior to OpenGraph.
  • -Anc: This version incorporates only sequence sampling. It significantly reduces memory costs during the training phase. Additionally, by focusing on the current training context, it achieves the best performance on OGBL-ddi.
  • -Seq: This model removes token sequence sampling from OpenGraph, resulting in a significant decrease in training efficiency. However, the anchor sampling strategy in this version greatly reduces computational costs during the test phase. Despite the computational benefits, the anchor sampling strategy leads to a drop in performance compared to the full-version OpenGraph.
Table 3: Impact of sampling strategies on the efficiency and performance in the scalable graph transformer.
OGBL-ddi Memory Time R@20
Train Test Train Test
-S-A 5420MiB 1456MiB 22.72s 13.88s 0.0966
@@ 4.5 Impact of Sampling in Transformer (RQ4)
Ours 2358MiB 1202MiB 15.45s 12.09s 0.1006
ML-10M Memory Time R@20
Train Test Train Test
-S-A OOM OOM
@@
Ours 4470MiB 4550MiB 68.79s 54.17s 0.2816
4.6 Impact of Model Scale (RQ5)
@@
Hidden dimensionality. In contrast to the number of transformer layers, the hidden dimensionality leads to quadratic growth in computational time. This reflects the rapid increase in model capacity required to accommodate more complex structural data. Consequently, we observe significant improvements in model performance, surpassing the performance curve for the graph transformer layers. Despite the increased model capacity, the growth in hidden dimensionality facilitates the $\mathbb{R}^{N}\rightarrow\mathbb{R}^{d}$ projection, reducing information loss and enhancing the quality of the graph token sequence for the subsequent graph transformer layers.
@@
Figure 3: The impact of model scale on downstream performance and training/testing time (seconds).
4.7 End-to-End Training Performance (RQ6)
@@
5 Related Work
Graph Neural Networks have gained attention for their ability to model complex relations in graphs Wu et al. (2020); Chen et al. (2020). GNNs use message passing to propagate information from neighboring nodes Jin et al. (2021); Yuan et al. (2020). Representative methods include Graph Convolutional Networks (GCNs) Gao et al. (2018); Zhang et al. (2021) and Graph Attention Networks (GATs) Zhang et al. (2022b); Liao et al. (2019). OpenGraph is inspired by the Graph Transformer Yun et al. (2019); Hu et al. (2020), known for capturing global dependencies in graphs.
@@ 5 Related Work
Self-Supervised Graph Learning aims to address the limited labeled data issue in graph tasks. These methods leverage graph structure and patterns for data-efficient training Wu et al. (2021b); Lee et al. (2022); Xia et al. (2023); Xiao et al. (2022); Yang et al. (2024). Graph contrastive learning frameworks, such as GraphCL You et al. (2020) and SGL Wu et al. (2021a), create meaningful representations by contrasting positive and negative samples using stochastic data augmenters. Adaptive augmentation schemes like JOAO You et al. (2021) and GCA Zhu et al. (2021) have been proposed. DGCL Li et al. (2021) and UMGRL Mo et al. (2023) address disentangling factors in contrastive learning. However, these solutions struggle with generalization. OpenGraph enhances graph model generalization across different tasks.
@@
LLM-based Graph Analysis. Recent advancements in LLMs have prompted interest in utilizing them for enhanced graph comprehension and analysis Ren et al. (2024). GraphLLM Chai et al. (2023) and GraphQA Fatemi et al. (2023) transform graphs into natural language descriptions, enabling improved interpretation and reasoning with LLMs. Techniques like instruction tuning in GraphEdit Guo et al. (2024) and GraphGPT Tang et al. (2024) incorporate rich textual information from text-attributed graphs for fine-tuning LLMs. However, in certain domains like user behavior graphs and neuronal graphs, obtaining high-quality textual features associated with graph nodes can be challenging. Therefore, there is a need for a graph model that can capture universal structural patterns from graphs, even in the absence of textual data.
6 Conclusion
@@
This research aims to develop an adaptable framework for capturing complex topological patterns in diverse graph structures. Our model demonstrates exceptional generalization capabilities in zero-shot graph learning tasks across various applications. We utilize a scalable graph transformer architecture and LLM-enhanced data augmentation for efficiency and robustness. Extensive experiments on benchmark datasets validate our model’s performance. Future plans include incorporating counterfactual learning to discover noisy connections and influential structures while learning universal and transferable structural patterns in diverse graphs.
7 Limitations
Our study serves as an initial exploration of graph foundation models, focusing on distilling the generalization capabilities from LLMs without relying on textual features. However, it is crucial to acknowledge and address the limitations that require further attention in future studies.
Firstly, it is important to note that OpenGraph currently does not include modeling for heterogeneous relations and node types. This limitation may impact its performance and generalization capabilities, particularly when dealing with graph data that exhibits strong heterogeneity, such as knowledge graphs. Future research should prioritize the incorporation of heterogeneous representation learning modules to enable a wider range of applications.
Secondly, graphs are highly versatile data structures that can be applied to a wide range of domains. While we have evaluated the performance of OpenGraph on 8 datasets spanning different domains, it is crucial to further strengthen the experimental validation by testing it on additional datasets from even more diverse application domains. This will provide a more comprehensive understanding of its effectiveness and applicability.
Lastly, while OpenGraph shows promising generalization capability, its explainability remains unexplored. This not only hinders its applicability due to its black box nature, but also hinders us from gaining a deeper understanding of the underlying principles that drive its strong generalization capabilities. Future research should prioritize investigating techniques and methodologies to enhance the explainability of OpenGraph, allowing researchers and practitioners to gain insights into the internal workings of the model and to ensure its reliable and transparent deployment in real-world applications.
In summary, future studies should focus on incorporating heterogeneous representation learning, expanding the range of tested datasets, and enhancing the explainability of OpenGraph for broader applicability and deeper insights.